# Molecular Generative Adversarial Network with Multi-Property Optimization

### v1

@@ 3.3 Objective Function
In the iterative adversarial process, the generator is updated by the MCTS-based RL algorithm through the sampling of numerous samples, leading to a computational training process. In contrast, InstGAN utilizes rewards derived from the actor-critic RL algorithm and MIE for the calculation. The overall loss function for the generator is calculated as
@@
where $b_{t}$ denotes the baseline using the global moving-average rewards (Sutton & Barto, 2018), which is calculated using both the mean reward $\bar{R}$ across the current batch and the previous baseline $b_{t-1}$. The calculation is as follows:
@@
where $V$ represents the size of chemical vocabulary.
Algorithm 1 outlines the training procedure for InstGAN with multi-property optimization. The generator, discriminator, and multiple chemical property prediction networks are first pretrained. Then, the generator is trained to generate a dataset. Afterward, the discriminator and critics are updated using both real and generated SMILES strings. Finally, the IR and GR are calculated to update the generator.
@@ 4.1 Experimental Setup
Table 1: Statistical descriptions of ZINC and ChEMBL datasets.
@@
Datasets.
The performance of InstGAN was validated through experiments on two widely used chemical datasets: ZINC (Irwin et al., 2012) and ChEMBL (Gaulton et al., 2017). The ZINC dataset contains $250,000$ drug-like molecules. The ChEMBL dataset includes approximately 1.6 million molecules, with each having a median of 27 and a maximum of 88 heavy atoms.
Chemical properties.
The molecular property optimization focused on the following chemical properties:
  • Drug-likeness (QED) is a metric that quantifies the likelihood of a molecule being a successful drug based on favorable molecular characteristics (Bickerton et al., 2012).
  • Solubility (logP) measures how well a molecule dissolves in lipid versus aqueous environments, quantified as the Log of the partition coefficient (Comer & Tam, 2001).
  • Synthesizability (SA) is defined by the synthetic accessibility score, evaluating the ease with which a molecule can be synthesized (Ertl & Schuffenhauer, 2009).
  • Dopamine receptor D2 (DRD2) is a central nervous system G protein-coupled receptor that is essential for dopamine-mediated signaling (Olivecrona et al., 2017).
Metrics.
The assessment of the molecular generation follows criteria outlined in (Li & Yamanishi, 2023). Validity is assessed by the proportion of chemically valid molecules among all generated ones, validated practically using the RDKit tool (Landrum, 2013). Uniqueness is determined by the proportion of non-duplicated molecules among all valid ones. Novelty is defined as the proportion of unique molecules not present in the training set. Total is the ratio of novel molecules to all generated ones. Diversity is calculated as the average Tanimoto distance (Rogers & Tanimoto, 1960) between the Morgan fingerprints (Cereto-Massagué et al., 2015) of novel molecules (refer to Appendix B).
All these properties and evaluation metrics are normalized to a range of $[0,1]$, where a higher score indicates better performance. Table 1 provides statistics of the datasets.
Hyperparameters.
For pretraining, the learning rate was set to $42e-$ 4. InstGAN was pretrained with up to $60,000$ steps, the batch size was $1024$, the MIE trade-off $\beta$ was set to $1\mathrm{e}{-3}$, and the GR weight $W^{D}$ was $7.1\mathrm{e}{-5}$. For multi-property optimization, QED, logP, and SA were used as the desired chemical properties, and the weights of $W_{C_{i}}$ were set to 0.4, 0.3, and 0.3. The learning rate was set to $5\mathrm{e}{-6}$, and the GAN was trained with a maximum of $5,000$ steps with a batch size of $256$. Additionally, the dropout rate was 0.1, and the gradient clippings range was from -0.1 to 0.1. The MIE trade-off $\beta$ was set to $2.2\mathrm{e}{-2}$, the GR weight $W^{D}$ was $1\mathrm{e}{-5}$, and the smoothing factor $\alpha$ was set to 0.9. The L2 coefficient for the critics was $1\mathrm{e}{-6}$. The maximum string length of generated SMILES was fixed at 49. Evaluations were conducted every $100$ steps, with $10,000$ samples generated. All experiments were performed within the NVIDIA GeForce RTX 3090 GPU environment.
@@
VAE-based RNN-Attention (Dollar et al., 2021) 71.57 99.94 100.0 71.53
TransVAE (Dollar et al., 2021) 25.39 99.96 100.0 25.38
Character-VAE (Kusner et al., 2017) 86.65 81.21 26.36 18.55
Grammar-VAE (Kusner et al., 2017) 91.91 77.24 11.90 8.45
JT-VAE (Jin et al., 2018) 100.0 19.75 99.75 19.70
Flow-based GraphAF (Shi et al., 2020) 68.00 99.10 100.0 67.39
GraphDF (Luo et al., 2021) 89.03 99.16 100.0 88.28
MoFlow (Zang & Wang, 2020) 81.76 99.99 100.0 81.75
GraphCNF (Lippe & Gavves, 2021) 63.56 100.0 100.0 63.56
Diffusion-based GDSS (Jo et al., 2022) 97.01 99.64 100.0 96.66
D2L-OMP (Guo et al., 2023) 97.51 99.88 100.0 97.39
GAN-based ORGAN (Guimaraes et al., 2017) 67.96 98.20 98.39 65.66
MolGAN (De Cao & Kipf, 2018) 95.30 4.30 100.0 4.10
TransORGAN (Li et al., 2022) 74.31 91.79 100.0 68.21
SpotGAN (Li & Yamanishi, 2023) 93.26 92.78 92.75 80.25
InstGAN Pretrain (Average) 95.45 98.63 99.71 93.87
Single property (QED) 97.89 98.31 99.69 95.94
Single property (logP) 96.65 98.42 99.93 95.05
Single property (SA) 97.46 98.59 99.75 95.85
Multi-properties (QED, logP, SA) 97.71 98.71 99.64 96.10
@@ 4.2 Evaluation Results
Table 2 compared the results of InstGAN with various baseline models (including VAE-, flow-, diffusion-, and GAN-based models) for chemical property optimization on the ZINC dataset. To ensure a fair comparison, InstGAN was pretrained several times, and the average results are presented. Additional details on these multiple pretraining sessions are provided in Appendix C. For VAE-based models, although RNN-Attention and TransVAE generated molecules that were highly unique and novel, their validity (i.e., $<71.6\%$) was significantly lower compared to InstGAN. InstGAN outperformed Character-VAE and Grammar-VAE in all metrics. Although JT-VAE exhibited high validity and novelty, the uniqueness was only $19.75\%$, significantly lower than that of InstGAN. For flow-based models, despite exhibiting high uniqueness and novelty, their validity was lower, specifically less than $89.03\%$, which significantly trails behind InstGAN. InstGAN demonstrated comparable performance to SOTA diffusion models GDSS and D2L-OMP. All models, including pretraining and tasks involving single properties (QED, logP, SA), as well as multi-property tasks (QED, logP, SA), achieved an overall score surpassing 93.87%. Among GAN-based models, while MolGAN and TransORGAN exhibited a novelty of 100.0%, their validity and uniqueness fell significantly lower than InstGAN. Specifically, MolGAN demonstrated a mere 4.3% uniqueness, and TransORGAN exhibited a validity of only 74.31%. ORGAN’s validity stood at 67.96%, markedly lower than InstGAN’s impressively high validity exceeding 95.45%. Furthermore, InstGAN outperformed SpotGAN in terms of validity, uniqueness, novelty, and total score, with the highest overall performance. InstGAN was trained for single- and multi-property optimization. In single-property optimization, InstGAN was trained individually using QED, logP, and SA. In multi-property optimization, these three properties were used to jointly train InstGAN. The validity, uniqueness, novelty, and total score all reached up to $93.87\%$. Overall, while InstGAN generated molecules from less informative SMILES strings compared to graph-based models, it surpassed VAE-, flow-, and GAN-based baselines and demonstrated comparable performance to SOTA diffusion models. This showcases InstGAN’s robust capability in molecular generation, excelling in both single- and multi-property optimization.
Table 3: Multi-property assessment and comparison of the top-k generated molecules.
Model (Property) Top-1 Top-1000 Top-2000 Top-4000 Top-6000 Top-8000
GraphAF (QED) 0.94 0.57 - - - -
GDSS (QED) 0.94 0.85 - - - -
MoFlow (QED) 0.93 0.78 - - - -
D2L-OMP (QED) 0.95 0.85 - - - -
InstGAN (QED) 0.95 0.93 0.92 0.90 0.88 0.86
InstGAN (logP) 1.00 0.89 0.85 0.81 0.78 0.74
InstGAN (SA) 1.00 0.99 0.98 0.96 0.94 0.92
@@
Table D.1 and Table 3 show the top-k property scores for the single- and multi-property optimization of the generated molecules, respectively. In single-property optimization, InstGAN enhanced all targeted chemical properties. The QED score showed a 30.1% improvement (from 0.73 to 0.95) for Top-1 and a 21.9% improvement (from 0.73 to 0.89) for Top-8000, compared with the training dataset. Furthermore, the generated molecules of InstGAN with logP and SA as the desired properties exhibited a notable improvement, with logP scores increasing by 78.6% for Top-1 and 76.8%, as well as 67.9% for SA scores for Top-8000, respectively. In multi-property optimization, InstGAN generated molecules with higher QED scores, compared with other baselines. Especially, InstGAN improved the QED score of the Top-1000 by 9.4%, comparing to the SOTA D2L-OMP baseline. Additionally, the generated molecules of InstGAN demonstrated substantial improvements, with the scores increasing by 30.1%, 78.6%, and 78.6% for Top-1, and 17.8%, 32.1%, and 64.3% for Top-8000, compared with the training dataset. Overall, InstGAN showcased substantial enhancements in both single- and multi-property optimization, underscoring its effectiveness in improving targeted chemical properties.
Figures D.1 and D.2 in the appendix show the top-ranked (Top-1) molecular structures generated by InstGAN in the single- and multi-property optimization tasks, respectively. The generated molecule adhered to Hückel’s rules (Klein & Trinajstic, 1984), essential for obtaining the targeted chemical properties of new drugs. These findings suggest that InstGAN successfully produced new drug-like molecules with relatively high QED, logP, and SA scores.
Figure 2: Trends in chemical properties of molecules generated during single-property optimization training.
Figure 2 and Fig. D.3 in the appendix depict the curves of average chemical property values versus training steps for molecules generated in single- and multi-property optimization tasks. Chemical properties for both single- and multi-property optimization exhibited gradual increases over 5000 training steps. In single-property optimization, the independence of the three chemical properties resulted in distinct and noticeable score increases. However, during the multi-property optimization process, the mutual constraints between properties led to similar changing trends in the values of the three properties.
(a) Drug-likeness (QED) distributions.
(b) Solubility (logP) distributions.
(c) Synthesizability (SA) distributions.
@@
Figure D.4 and Fig. 3 show the property distributions of molecules generated with single- and multi-property optimization, respectively. Intuitively, in comparison to the molecule distributions in the training set (in blue), the chemical property distributions of the newly generated molecules (in green) shifted to the right overall. This suggests that InstGAN produced a greater number of new molecules with desirable properties. Furthermore, the property scores of molecules generated through multi-property optimization were marginally lower than those from single-property optimization. This difference arises because InstGAN had to concurrently consider the enhancement of three properties (i.e., QED, logP, and SA) during multi-property optimization. In essence, multi-property optimization for molecular generation proves to be more challenging and intricate. Thus, the results demonstrated the effectiveness of InstGAN in desired chemical property optimization.
@@ 4.3 Ablation Studies
Table 4 demonstrates the impact of various InstGAN variants on molecular generation performance. During the training process of InstGAN, we excluded GR, and MIE individually to create three distinct variants, namely, “w/o IR,” “w/o GR,” and “w/o MIE.” These variants were then compared with InstGAN. In the “w/o IR” scenario, the calculation of instant reward was substituted with the MCTS RL algorithm. While MCTS enhanced GAN training stability with its extensive sampling, resulting in relatively high validity (97.80%), its computational complexity, stemming from the large number of samples, constrains its applicability in lengthy sequences. In the case of “w/o GR”, the validity was the lowest (95.96%, compared to 97.56% of InstGAN). Given that the GR involves the global information of a SMILES string, it provides additional sequence-related information for generating subsequent tokens in the molecular auto-regression process, thereby contributing to the validity improvement. MIE, by smoothing the probabilities of generating tokens, enhances diversity in sampling tokens with non-maximum probabilities. Consequently, in the “w/o MIE” scenario, the generated molecular distribution exhibited the lowest uniqueness (96.50%, compared to 98.47% of InstGAN). InstGAN, incorporating IR, GR, and MIE, showcased high validity, uniqueness, and Novelty scores, and achieved the highest total score of 95.81%.
Additionally, Tables E.1, E.2, E.3, and E.4 display the effects of $\lambda$ on the performance. Overall, InstGAN can be applied to extensive chemical databases, optimizing chemical properties while retaining a low computational cost.
Table 4: Effect of different variants of InstGAN.
@@ 4.4 Case Studies
Table F.1 in Appendix F assesses the performance of QED and DRD2 properties, demonstrating that the QED and DRD2 scores change with the corresponding weights. Furthermore, InstGAN enhanced the bioactivity of the generated molecules to 97.21%. Additionally, we selected a QED and DRD2 weight ratio of (0.3, 0.7) and generated bioactive molecules. Figure 4 compares the generated molecules with high DRD2 scores to similar approved drugs. These Top-3 molecules generated by InstGAN have high QED and DRD2 scores and are highly similar to approved drugs in the ChEMBL database, proving the effectiveness of InstGAN.

### v2

@@ 3.3 Objective Function
In the iterative adversarial process, the generator is updated by the MCTS-based RL algorithm through the sampling of numerous samples, leading to a computational training process. In contrast, InstGAN utilizes rewards derived from the actor-critic RL algorithm and MIE for the calculation. Please see the description of the MCTS-based RL and actor-critic RL for algorithms comparison in Appendix B. The overall loss function for the generator is calculated as
@@
where $b_{t}$ denotes the baseline using the global moving-average rewards Sutton and Barto (2018), which is calculated using both the mean reward $\bar{R}$ across the current batch and the previous baseline $b_{t-1}$, calculated as
@@
where $V$ represents the size of chemical vocabulary. Algorithm 1 outlines the training procedure for InstGAN. The generator, discriminator, and multiple chemical property prediction networks are first pre-trained. Then, the generator is trained to generate a dataset. Afterward, the discriminator and critics are updated using both real and generated SMILES strings. Finally, the IR and GR values are calculated to update the generator’s parameters.
@@ 4.1 Experimental Setup
Datasets.
The performance of InstGAN was validated through experiments on two widely used chemical datasets: ZINC Irwin et al. (2012) and ChEMBL Gaulton et al. (2017). The ZINC dataset contains $250,000$ drug-like molecules. The ChEMBL dataset includes approximately 1.6 million molecules, with each having a median of 27 and a maximum of 88 heavy atoms.
Chemical properties.
Drug-likeness (QED) quantifies the probability that a molecule belongs to a drug Bickerton et al. (2012). Solubility (logP) measures how well a molecule dissolves in lipid versus aqueous environments, quantified as the Log of the partition coefficient Comer and Tam (2001). Synthesizability (SA) is defined by the synthetic accessibility score, evaluating the ease with which a molecule can be synthesized Ertl and Schuffenhauer (2009). Dopamine receptor D2 (DRD2) is a central nervous system G protein-coupled receptor that is essential for dopamine-mediated signaling Olivecrona et al. (2017). Table 1 provides detailed statistical descriptions of the datasets.
Metrics.
@@
Table 1: Statistical descriptions of the datasets.
Validity is assessed by the proportion of chemically valid molecules among all generated ones, validated practically using the RDKit tool Landrum (2013). Uniqueness is determined by the proportion of non-duplicated molecules among all valid ones. Novelty is defined as the proportion of unique molecules not present in the training set. Total is the ratio of novel molecules to all generated ones. Diversity is calculated as the average Tanimoto distance Rogers and Tanimoto (1960) between the Morgan fingerprints Cereto-Massagué et al. (2015) of novel molecules. All these properties and metrics are normalized to a range of $[0,1]$, with a higher score indicating better performance. For a detailed description of the chemical properties, evaluation metrics and hyperparameters, please refer to Appendix C.
@@
VAE-based
RNN-Attention Dollar et al. (2021) 71.57 99.94 100.0 71.53
TransVAE Dollar et al. (2021) 25.39 99.96 100.0 25.38
Character-VAE Kusner et al. (2017) 86.65 81.21 26.36 18.55
Grammar-VAE Kusner et al. (2017) 91.91 77.24 11.90 8.45
JT-VAE Jin et al. (2018) 100.0 19.75 99.75 19.70
Flow-based
GraphAF Shi et al. (2020) 68.00 99.10 100.0 67.39
GraphDF Luo et al. (2021) 89.03 99.16 100.0 88.28
MoFlow Zang and Wang (2020) 81.76 99.99 100.0 81.75
GraphCNF Lippe and Gavves (2021) 63.56 100.0 100.0 63.56
Diffusion-based
GDSS Jo et al. (2022) 97.01 99.64 100.0 96.66
D2L-OMP Guo et al. (2023) 97.51 99.88 100.0 97.39
GAN-based
ORGAN Guimaraes et al. (2017) 67.96 98.20 98.39 65.66
MolGAN De Cao and Kipf (2018) 95.30 4.30 100.0 4.10
TransORGAN Li et al. (2022) 74.31 91.79 100.0 68.21
SpotGAN Li and Yamanishi (2023) 93.26 92.78 92.75 80.25
InstGAN
Pre-train (Average) 95.45 98.63 99.71 93.87
Property (QED) 97.89 98.31 99.69 95.94
Property (logP) 96.65 98.42 99.93 95.05
Property (SA) 97.46 98.59 99.75 95.85
Multi-property 97.71 98.71 99.64 96.10
@@ 4.2 Evaluation Results
Table 2 compared the results of InstGAN with various baseline models (including VAE-, flow-, diffusion-, and GAN-based models) for chemical property optimization on the ZINC dataset. To ensure a fair comparison, InstGAN was pre-trained several times, and the average results are presented. Additional details on these multiple pre-training sessions are provided in Appendix D. For VAE-based models, although RNN-Attention and TransVAE generated molecules that were highly unique and novel, their validity (i.e., $71.6\%$) was significantly lower compared to InstGAN. InstGAN outperformed Character-VAE and Grammar-VAE in all metrics. Although JT-VAE exhibited high validity and novelty, the uniqueness was only $19.75\%$, significantly lower than that of InstGAN. For flow-based models, despite exhibiting high uniqueness and novelty, their validity was lower, specifically less than $89.03\%$, which significantly trails behind InstGAN. InstGAN demonstrated comparable performance to SOTA diffusion models GDSS and D2L-OMP. All models, including pre-training and tasks involving single properties (QED, logP, SA), as well as multi-property tasks (QED, logP, SA), achieved an overall score surpassing 93.87%. Among GAN-based models, while MolGAN and TransORGAN exhibited a novelty of 100.0%, their validity and uniqueness fell significantly lower than InstGAN. Specifically, MolGAN demonstrated a mere 4.3% uniqueness, and TransORGAN exhibited a validity of only 74.31%. ORGAN’s validity stood at 67.96%, markedly lower than InstGAN’s impressively high validity exceeding 95.45%. Furthermore, InstGAN outperformed SpotGAN in terms of validity, uniqueness, novelty, and total score, with the highest overall performance. InstGAN was trained for single- and multi-property optimization. In single-property optimization, InstGAN was trained individually using QED, logP, and SA. In multi-property optimization, these three properties were used to jointly train InstGAN. The validity, uniqueness, novelty, and total score all reached up to $93.87\%$. Overall, while InstGAN generated molecules from less informative SMILES strings compared to graph-based models, it surpassed VAE-, flow-, and GAN-based baselines and demonstrated comparable performance to SOTA diffusion models. This highlights InstGAN’s robust capability in molecular generation, demonstrating excellence in both single-property and multi-property optimization.
@@
Model (Property) Top-1 Top-5 Top-10 Top-100 Top-1000
GraphAF 0.94 0.93 0.92 0.86 0.57
GDSS 0.94 0.94 0.93 0.91 0.85
MoFlow 0.93 0.92 0.92 0.87 0.78
D2L-OMP 0.95 0.94 0.94 0.91 0.85
InstGAN 0.95 0.95 0.95 0.95 0.94
InstGAN (Multi) 0.95 0.95 0.95 0.94 0.93
Table 3: QED assessment of the top-k generated molecules.
Table 3 and Table E.1 show the top-k property scores for the multi- and single-property optimization of the generated molecules, respectively. In multi-property optimization, InstGAN enhanced all targeted chemical properties. The QED scores showed a 30.1% improvement (from 0.73 to 0.95) for Top-1 and 27.4% improvement (from 0.73 to 0.93) for Top-1000, compared with the training dataset. Furthermore, the generated molecules of InstGAN with logP and SA as the desired properties exhibited a notable improvement, with logP scores increasing by 78.6% for Top-1 and 58.9%, as well as 76.8% for SA scores for Top-1000, respectively. In both property optimization, InstGAN generated molecules with higher QED scores, compared with other baselines. Especially, InstGAN improved the QED score of the Top-1000 by 9.4% and 10.6%, comparing to the SOTA D2L-OMP baseline. In single-property optimization, the generated molecules of InstGAN demonstrated substantial improvements, with the scores increasing by 30.1%, 78.6%, and 78.6% for Top-1, and 28.8%, 78.6%, and 78.6% for Top-1000, compared with the training dataset. Overall, InstGAN showcased substantial enhancements in property optimization, underscoring its effectiveness in improving targeted chemical properties.
Figures E.1 and E.2 in the appendix show the top-ranked (Top-1) molecular structures generated by InstGAN in the single- and multi-property optimization tasks, respectively. The generated molecule adhered to Hückel’s rules Klein and Trinajstic (1984), essential for obtaining the targeted chemical properties of new drugs. These findings suggest that InstGAN successfully produced new drug-like molecules with relatively high QED, logP, and SA scores.
Figures E.3 and E.4 in the appendix depict the curves of average chemical property values versus training steps for molecules generated in single- and multi-property optimization tasks. Chemical properties for both single- and multi-property optimization exhibited gradual increases over 5000 training steps. In single-property optimization, the independence of the three chemical properties resulted in distinct and noticeable score increases. Moreover, during the multi-property optimization process, the mutual constraints between properties led to similar changing trends in the values of the three properties.
@@
Fig. 2 and Figure E.5 in the Appendix show the property distributions of molecules generated with multi- and single-property optimization, respectively. Intuitively, in comparison to the molecule distributions in the training set (in blue), the chemical property distributions of the newly generated molecules (in green) shifted to the right overall. This suggests that InstGAN produced a greater number of new molecules with desirable properties. Furthermore, the property scores of molecules generated through multi-property optimization were marginally lower than those from single-property optimization.
This difference arises primarily because InstGAN had to concurrently consider the enhancement of three properties and several of which are inherently conflicting during multi-property optimization. Detailed results and trade-off analyses are provided in Appendix F. In essence, multi-property optimization for molecular generation proves to be more challenging and intricate.The results demonstrated the effectiveness of InstGAN in chemical property optimization.
@@ 4.3 Ablation Studies
Table 4 demonstrates the impact of various InstGAN variants on molecular generation performance. During the training process of InstGAN, we excluded GR, and MIE individually to create three distinct variants, namely, “w/o IR,” “w/o GR,” and “w/o MIE.” These variants were then compared with InstGAN. In the “w/o IR” scenario, the calculation of instant reward was substituted with the MCTS-based RL algorithm. While MCTS enhanced GAN training stability with its extensive sampling,
@@
Table 4: Effect of different variants of InstGAN.
resulting in relatively high validity (97.80%), its computational complexity, stemming from the large number of samples, constrains its applicability in lengthy sequences. In the case of “w/o GR”, the validity was the lowest (95.96%, compared to 97.56% of InstGAN). Given that the GR involves the global information of a SMILES string, it provides additional sequence-related information for generating subsequent tokens in the molecular auto-regression process, thereby contributing to the validity improvement. MIE, by smoothing the probabilities of generating tokens, enhances diversity in sampling tokens with non-maximum probabilities. Consequently, in the “w/o MIE” scenario, the generated molecular distribution exhibited the lowest uniqueness (96.50%, compared to 98.47% of InstGAN). InstGAN, incorporating IR, GR, and MIE, achieved the highest total score of 95.81%.
Additionally, Tables G.1, G.2, G.3, and G.4 display the effects of $\lambda$ on the performance. InstGAN can be applied to extensive chemical databases, optimizing chemical properties while retaining a low computational cost.
Integrating an actor-critic RL with instant and global rewards is crucial to the success of InstGAN. When compared to MCTS sampling, our strategy is more computationally eifficient, and its fine-grained feedback over SMILES sub-strings provides informative gradients at the early training stage mitgating the gradient vanishing problem that often hinders discrete GANs. The same mechanism also stabilizes the learning process: in the integration setup of GAN and RL, the generator simultaneously minimizes KL divergence and maximizes JS divergence, producing conflicting gradients Saxena and Cao (2021). Injecting entropy through token-level rewards and MIE can balance the two objectives and aligns with RL findings that entropy regularization facilitates exploration and robust learning Ahmed et al. (2019). As a result, InstGAN achieves more stable convergence, improved sample quality, and performance on par with SOTA models.
@@ 4.4 Case Studies
Table H.1 in Appendix H assesses the performance of QED and DRD2 properties, demonstrating that the QED and DRD2 scores change with the corresponding weights. Furthermore, InstGAN enhanced the bioactivity of the generated molecules to 97.21%. Additionally, we selected a QED and DRD2 weight ratio of (0.3, 0.7) and generated bioactive molecules. Figure 3 compares the generated molecules with high DRD2 scores to similar approved drugs. These Top-3 molecules generated by InstGAN have high QED and DRD2 scores and are highly similar to approved drugs in the ChEMBL database, proving the effectiveness of InstGAN.